OpenStudy (anonymous):

The initial term and common difference are given. What is the nth term of the arithmetic sequence? a1 = -2; d = 4

OpenStudy (anonymous):

Well each term differs from the previous one by 4. So a2 = a1 + 4 a3 = a2 + 4 a4 = a3 + 4 etc.

OpenStudy (anonymous):

so 4 + 4n?

OpenStudy (anonymous):

I disagree that it should be 4 + 4n. Probably should be a(n) = -6 + 4n Since then when n = 1, You'll have a(1) = -6 + 4 = -2 And a1 was given to be -2 so that's right.